if k multiply velocity multiply mass is displacement, then the displacement of k are
1
Expert's answer
2020-10-29T07:01:53-0400

Solution

k multiply velocity multiply mass is displacement

Then expressed as

"k\times \frac{meter}{sec}\times Kg=meter"

"k=\frac{sec}{Kg}"

This is dimension of k.